Suppose a bank enters a repurchase agreement in which it agrees to buy Treasury securities from a correspondent bank at a price of $410 million, with the promise to sell them back at a price of $410.1 million in five days. Calculate the quoted and bond equivalent yield. What is the actual yield if you consider reinvestment and compounding?
Answer #1

Bank Buys at a PURCHASE PRICE of $ 410 million

Sells the Investment at a SALE PRICE of  $ 410.10 million

Time of Holding is 5 Days

Profit on Value Terms = Sale Price - Purchase Price = $ (410.10 - 410 ) million = 0.10 million

Absolute Yield = Absolute Yield / Purchase Price *100

= 0.10 / 410 * 100 = 0.0244 %

Yield in Percentage on the basis of Whole Year Investment = Absolute Yield / Purchase Price / No. of Days of Holding * 365 ( whole Year ) * 100

= 0.10 / 410 / 5 *365 * 100

= 1.7805 %

THIS IS BASED ON THE ABOVE INFORMATION PROVIDED.